OpenStudy (anonymous):

Write the product as a trinomial. (m + 2)(m – 7)

OpenStudy (anonymous):

I'm confused please help

OpenStudy (dtan5457):

you just use foil m times m=m^2 m times -7 =-7m m times 2=2m -7 times 2=-14 that equates to m^2-7m+2m-14 aka m^2-5m-14